# The Influence of V-Defects, Leakage, and Random Alloy Fluctuations on the Carrier Transport in Red InGaN MQW LEDs

### v1

@@ Acknowledgements.
This project is supported by National Science and Technology Council (NSTC) under grant No. 112-2221-E-002-215-MY3, 113-2640-E-002-005, and 113-2124-M-002-013-MY3. Support at UCSB was provided by the Solid State Lighting and Energy Electronics Center (SSLEEC); U.S. Department of Energy under the Office of Energy Efficiency & Renewable Energy (EERE) Award No. DE-EE0009691; the National Science Foundation (NSF) RAISE program (Grant No. DMS-1839077); the Simons Foundation (Grant #s 601952 and 1027114 for JSS and CW, respectively).
@@
Yuh-Renn Wu conceptualization (equal); formal analysis (equal); funding acquisition (lead); project administration (lead); software(lead); supervision (lead); writing-review and editing (lead).

### v2

@@ IV Conclusion
V SUPPLEMENTARY MATERIAL
Supplementary material associated with this article includes: (1) sensitivity analyses of simulated IQE to non-radiative lifetimes (Figs. S1-S4); (2) methodology for indium composition determination (Fig. S5) and its impact on emission (Fig. S6); and (3) a detailed simulation execution guide with input files. All supplementary files are available online at [URL will be inserted by publisher].
@@ Acknowledgements.
This project is supported by National Science and Technology Council (NSTC) under grant No. 112-2221-E-002-215-MY3, 112-2221-E-002-214-MY3, 113-2640-E-002-005, and 113-2124-M-002-013-MY3 and by Ministry of Education UAAT-KOOU project (113M7056). Support at UCSB was provided by the Solid State Lighting and Energy Electronics Center (SSLEEC); U.S. Department of Energy under the Office of Energy Efficiency & Renewable Energy (EERE) Award No. DE-EE0009691; the National Science Foundation (NSF) RAISE program (Grant No. DMS-1839077); the Simons Foundation (Grant #s 601952 and 1027114 for JSS and CW, respectively).
@@
Yuh-Renn Wu: conceptualization (equal); formal analysis (equal); funding acquisition (lead); project administration (lead); software(lead); supervision (lead); writing-review and editing (lead).